# Exposing Barriers to Flexibility Aggregation in Unbalanced Distribution Networks

### v2

@@ III-A Case Study: Illustrative 5-bus System
Unbalanced network, $\phi=\{a\}$ 177.03 (66.07%)
Unbalanced network, $\phi=\{b\}$ 165.86 (61.90%)
Unbalanced network, $\phi=\{c\}$ 166.16 (62.01%)
@@
However, for existing distribution systems, the perfect coordination assumption may be unreasonable since 1) single-phase flexible units can be connected at different locations and belong to different owners, and 2) required automation and communication technologies are not widely adopted. To analyse the lack of unit coordination, phase coordination constraint (1p) is imposed. Under this constraint, flexibility service in phase $\phi$ can be provided only by the single-phase unit connected to this phase. Other units cannot manage the voltage unbalance or assist in maximising network flexibility. The impact of phase coordination on P-Q flexibility areas is shown in Fig. 6, where VUF constraints (1o) in the range 0.1-1.0% are iteratively imposed. The findings of this simulation are twofold: (i) without coordination between units connected to different phases, the limits of aggregated flexibility reduce significantly, for example, 60.76% flexibility area reduction for phase $\phi=\{b\}$; (ii) voltage unbalance limits have a major impact on flexibility services provided by units with a lack of coordination. It follows that a lack of phase coordination combined with voltage unbalance limits creates the worst conditions for providing flexibility services: single-phase units cannot effectively manage voltage unbalance across the network, which results in the infeasibility of services provision.
Figure 6: Analysis of the phase coordination and voltage unbalance impacts on flexibility services in the 5-bus system: (a) for phase $\phi=\{a\}$; (b) for phase $\phi=\{b\}$; (c) for phase $\phi=\{c\}$. The cross markers correspond to the initial operating point (with no flexibility provision), while the coordinates represent the total network’s power consumption. The services are provided by three single-phase flexible units. Phase coordination constraints imposed do not allow units connected to different phases to jointly provide flexibility and manage voltage unbalance, thus reducing flexibility services for each phase.
a)Refer to caption  b)Refer to caption  c)Refer to caption
The findings illustrated in this subsection emphasise the need for considering phase unbalance impacts on flexibility services in distribution networks. If ignoring these impacts, the potential of DER flexibility services may be greatly overestimated. The following subsection presents a realistic case study to demonstrate the scalability of the proposed framework.
III-B Case Study: 221-bus LV Distribution Network in the UK
The proposed flexibility quantification framework is applied to a real anonymised 0.4 kV distribution network located in the North West of England [39]. The network is visualised in Fig. 7 as a graph using the force-directed graph layout algorithm ForceAtlas2 [40]. The figure illustrates the network topology and the location of power demands across the network, where circles of different sizes correspond to the total apparent power consumed at each bus. The smallest circles indicate buses with no loads, while the largest circles stand for buses with a load of 3.52 kVA. The distribution of loads across phases is given in Table II. The network is supplied from a single three-phase balanced source, the primary substation (displayed as a triangle in Fig. 7). Yet, due to the uneven distribution of single-phase loads, the network experiences a voltage unbalance, with the highest VUF of 0.78% over all buses. Such a low level of voltage unbalance is considered acceptable according to existing grid codes [35, 36]. However, this voltage unbalance can still impact the network’s flexibility and limit the operation of DER.
@@ III-B Case Study: 221-bus LV Distribution Network in the UK
To investigate the impact of phase unbalance on flexibility services, 12 single-phase flexible units (four units per phase) were placed in the network as shown in Fig. 7. Each unit has a P-Q capability of ±5 kW and ±5 kVAr. That is, units can independently regulate their active and reactive power output in the range of [-5.0,5.0] kW and [-5.0,5.0] kVAr. Then, the aggregated P-Q flexibility areas were calculated per each phase by iteratively solving the flexibility estimation model (1a)-(1p). The impact of phase unbalance and coordination on P-Q flexibility areas is shown in Fig. 8, where VUF constraints (1o) in the range 0.5-1.0% are imposed. [^5] To avoid scalability issues, in the 221-bus system, VUF constraints (1l)-(1o) are imposed only on the 7 buses with the highest voltage unbalances.
Similar to the previous analysis for the simple 5-bus system, the P-Q areas in Fig. 8 indicate that the network’s flexibility is maximal when all single-phase flexible units are perfectly coordinated (as shown by the dashed lines). However, the absence of coordination between units connected to different phases leads to significant losses in the aggregated DER flexibility potential. For example, the P-Q flexibility area for phase $\phi=\{a\}$ reduces by 30% (from 1721.7 to 1211.6 kVA2). Moreover, imposing voltage unbalance constraints (1o) further limits the operation of flexible units and significantly reduces the network’s aggregated flexibility.
@@
Figure 8: Analysis of the phase coordination and voltage unbalance impacts on flexibility services in the real 221-bus system: (a) for phase $\phi=\{a\}$; (b) for phase $\phi=\{b\}$; (c) for phase $\phi=\{c\}$. The cross markers correspond to the initial operating point (with no flexibility provision), while the coordinates represent the total network’s power consumption. The services are provided by 12 single-phase flexible units. Phase coordination constraints imposed do not allow units connected to different phases to jointly provide flexibility and manage voltage unbalance, thus reducing flexibility services for each phase.
a)Refer to caption  b)Refer to caption  c)Refer to caption
@@
Note that in this simulation setting, all flexible units have been assigned the same P-Q capabilities and distributed evenly across phases. Yet, their operation is limited considerably due to phase coordination issues and voltage unbalance. It follows that the provision of flexibility services by DER can be challenging even in LV networks with low load unbalance as flexible units can exacerbate existing voltage unbalances by producing or consuming power. To effectively manage such unbalances and voltage constraints, units connected to different phases must be perfectly coordinated. The simulations performed for the realistic 221-bus network demonstrate that lack of coordination between units significantly reduces the potential of DER flexibility services and can magnify voltage unbalance problems. Other aspects and implications of these findings are discussed in Section IV.
III-C Stochastic analysis
@@ IV Discussion
The simulations performed in Section III demonstrate the impacts of phase unbalance and imperfect coordination between distributed flexibility providers. Specifically, it is shown that the total DER flexibility potential can be significantly reduced due to voltage unbalance constraints and lack of coordination between flexible units connected to different phases. This highlights the importance of accurate nonlinear three-phase OPF models for the analysis of the network’s aggregated P-Q flexibility, as such models are capable of simulating voltage unbalances and coordination issues across phases. In this regard, the proposed flexibility quantification framework allows to directly translate voltage unbalance and phase coordination constraints as reductions of the aggregated P-Q flexibility. Thus, it can be used by DSO to realistically predict flexibility available in distribution networks and avoid severe voltage unbalances.
@@
Multiple aspects of flexibility provision in unbalanced LV distribution networks have to be further investigated. One of the key factors affecting the aggregated P-Q flexibility is the relationship between the placement of flexible units and locations of voltage unbalances. For instance, in the 221-bus case study displayed in Fig. 7, the highest load unbalance is happening at the feeder with the flexible unit F5{a} connected. Six more flexible units are placed in this part of the network. Therefore, the design of this experiment inherently creates constraints on the operation of these flexible units due to the nearby voltage unbalances. That is, units must be coordinated when consuming or producing power to manage voltage constraints and keep the network’s operation feasible. For example, units connected to phase $\phi=\{a\}$ (the phase with the highest load) must not increase their power consumption while units in other phases consume power. Otherwise, voltage unbalance in some parts of the network may increase to unacceptable levels or even the voltage limits (1b) for phase $\{a\}$ may be violated. Note that this would not happen if flexible units were located far from the load unbalance, say, at the primary substation. Therefore, the question arises: how does the initial network unbalance impact the network’s P-Q flexibility and how does it restrict flexible resources at different locations? This question has implications for DER planning models. Depending on DSO metrics and objectives, some DER should be placed closer to the primary substation, while other flexible resources should be located closer to existing voltage unbalances to manage voltage constraints. A similar problem was considered in [23, 24], where it was demonstrated that suboptimal selection of the DER connection phase can lead to voltage unbalance and limit DER capacity.
@@
Regarding the computational aspects of flexibility quantification, the main computational burden comes from the need to iteratively solve the nonlinear optimisation problem (1a)-(1p). A single OPF problem with no voltage unbalance constraints converges in 0.1-0.2 seconds and 5-50 seconds for the 5-bus and 221-bus cases, respectively. This means that the boundary of an aggregated P-Q flexibility area with 50 points can be estimated in less than 2500 seconds (43 minutes). On average, it takes 16 minutes to calculate a single boundary presented in Fig. 8. Note that these values are reported for the simulations performed on a laptop with an Intel Core i7-10510U CPU 1.80GHz and 16 GB of RAM. The flexibility estimation process can be greatly sped up by using more powerful computers and parallel computing (solving OPF problems for different points on different machines). Therefore, DSO can perform this flexibility assessment close to real time.

### v3

@@ IV-A Case Study: Illustrative 5-bus System
Unbalanced network, $\phi=\{a\}$ 177.03 (166.07 %)
Unbalanced network, $\phi=\{b\}$ 165.86 (161.90 %)
Unbalanced network, $\phi=\{c\}$ 166.16 (162.01 %)
@@
However, for existing distribution systems, the assumption of perfect coordination may be unrealistic because: 1) single-phase flexible units can be connected at different locations and belong to different owners, and 2) required automation and communication technologies are not widely adopted. To analyse the lack of unit coordination, phase coordination constraint (1p) is imposed. Under this constraint, flexibility service in phase $\phi$ can be provided only by single-phase units connected to that phase. Other units cannot manage the voltage unbalance or assist in maximising network flexibility. The impact of phase coordination and voltage unbalance constraints on P-Q flexibility areas is shown in Fig. 8. The red shaded areas correspond to simulations in which VUF constraints (1o) in the range 0.1-1.0% are iteratively imposed. The quantitative results for these simulations are summarised in Table III, where aggregated flexibility areas are reported in kVAr2. Relative reductions in flexibility are expressed in percent per phase, with 100% denoting the maximum flexibility under perfect coordination and no voltage unbalance constraints. The findings of these simulations are twofold: (i) without coordination between units connected to different phases, aggregated flexibility areas reduce significantly, by over 54%, depending on the phase; (ii) voltage unbalance limits have a major impact on flexibility services when coordination between units is absent. It follows that a lack of phase coordination combined with voltage unbalance limits creates the worst conditions for providing flexibility services: single-phase units cannot effectively manage voltage unbalance across the network, which can make flexibility provision infeasible.
@@
Figure 8: Impact of phase coordination and voltage unbalance constraints on flexibility services in the 5-bus system. The aggregated P-Q flexibility areas are shown for each phase. Cross markers indicate the initial operating point with no flexibility provision, while the coordinates represent the total network power consumption per phase. Flexibility services are provided by three single-phase units. Imposing phase coordination constraints prevents joint flexibility provision by units connected to different phases, thereby limiting voltage unbalance management and reducing the feasible area for each phase.
TABLE III: Impact of Phase Coordination and Voltage Unbalance Constraints on Aggregated P-Q Flexibility in the 5-Bus System
Simulation Aggregated P-Q flexibility area per phase, kVAr2 (%)
Phase $\phi=\{a\}$ Phase $\phi=\{b\}$ Phase $\phi=\{c\}$
Perfect coordination of units, no voltage unbalance constraints $306.24$  ( $100.00$ %) $306.04$  ( $100.00$ %) $305.60$  ( $100.00$ %)
No coordination of units, no voltage unbalance constraints $138.87$  ( $45.35$ %) $120.10$  ( $39.24$ %) $115.56$  ( $37.81$ %)
No coordination of units, VUF limit of 1.0% imposed $138.23$  ( $45.14$ %) $101.90$  ( $33.30$ %) $110.88$  ( $36.28$ %)
No coordination of units, VUF limit of 0.9% imposed $135.26$  ( $44.17$ %) $83.23$  ( $27.20$ %) $97.34$  ( $31.85$ %)
No coordination of units, VUF limit of 0.8% imposed $129.73$  ( $42.36$ %) $65.81$  ( $21.50$ %) $82.83$  ( $27.11$ %)
No coordination of units, VUF limit of 0.7% imposed $120.06$  ( $39.20$ %) $49.37$  ( $16.13$ %) $67.64$  ( $22.13$ %)
No coordination of units, VUF limit of 0.6% imposed $103.91$  ( $33.93$ %) $32.48$  ( $10.61$ %) $53.74$  ( $17.59$ %)
No coordination of units, VUF limit of 0.5% imposed $84.93$  ( $27.73$ %) $17.68$  ( $5.78$ %) $41.4$  ( $13.55$ %)
No coordination of units, VUF limit of 0.4% imposed $62.71$  ( $20.48$ %) $7.04$  ( $2.30$ %) $30.53$  ( $9.99$ %)
No coordination of units, VUF limit of 0.3% imposed $35.37$  ( $11.55$ %) $0.83$  ( $0.27$ %) $21.06$  ( $6.89$ %)
No coordination of units, VUF limit of 0.2% imposed $15.75$  ( $5.14$ %) $0.00$  ( $0.00$ %) $12.54$  ( $4.10$ %)
No coordination of units, VUF limit of 0.1% imposed $3.97$  ( $1.30$ %) $0.00$  ( $0.00$ %) $3.54$  ( $1.16$ %)
Note: All simulations assume an unbalanced network, with flexibility provided by single-phase units connected to different phases.
The findings of this subsection emphasise the importance of considering phase unbalance impacts when assessing flexibility services in distribution networks. If these impacts are ignored, the potential of DER flexibility services may be greatly overestimated. The following subsection presents a realistic case study to demonstrate the scalability of the proposed framework.
IV-B Case Study: 221-bus LV Distribution Network in the UK
The proposed flexibility quantification framework is applied to a real anonymised 0.4 kV distribution network located in the North West of England [churkin_Zenodo]. The network is visualised in Fig. 7 as a graph using the force-directed graph layout algorithm ForceAtlas2 [Jacomy2014]. The figure illustrates the network topology and the location of power demands across the network, where circles of different sizes correspond to the total apparent power consumed at each bus. The smallest circles indicate buses with no loads, while the largest circles stand for buses with a load of 3.52 kVA. The distribution of loads across phases is given in Table II. The network is supplied from a single three-phase balanced source, the primary substation (displayed as a triangle in Fig. 7). Yet, due to the uneven distribution of single-phase loads, the network experiences a voltage unbalance, with the highest VUF of 0.78% over all buses. Such a low level of voltage unbalance is considered acceptable according to existing grid codes [EnergyNetworksAssociation1990, NationalGrid2014]. However, this voltage unbalance can still impact the network’s flexibility and limit the operation of DER.
To investigate the impact of phase unbalance on flexibility services, 12 single-phase flexible units (four units per phase) were placed in the network as shown in Fig. 7. The locations of the units were selected to reflect realistic DER deployments in LV networks and to highlight the interrelation between unit operation and voltage unbalance. Specifically, the units were placed at nodes located towards the end of feeders, where voltage unbalance effects are more pronounced and flexibility injections are more likely to be constrained. Each unit has a P-Q capability of ±5 kW and ±5 kVAr. That is, units can independently regulate their active and reactive power output in the range of [-5.0,5.0] kW and [-5.0,5.0] kVAr. Then, the aggregated P-Q flexibility areas were calculated per each phase by iteratively solving the flexibility estimation model (1a)-(1p). The impact of phase coordination and voltage unbalance constraints on P-Q flexibility areas in the 221-bus system is shown in Fig. 9. The red shaded areas indicate simulations in which VUF constraints (1o) are imposed. [^5] To avoid scalability issues, in the 221-bus system, VUF constraints (1l)-(1o) are imposed only on the 7 buses with the highest voltage unbalances and close to the flexible units, as shown in Fig. 7 by the black cross markers. A detailed discussion of the voltage unbalance modelling and scalability properties is provided in Section V. The quantitative results are summarised in Table IV.
Similar to the previous analysis for the simple 5-bus system, the P-Q areas in Fig. 9 indicate that the network’s flexibility is maximised when all single-phase flexible units are perfectly coordinated, as shown by the dashed boundaries. However, the absence of coordination between units connected to different phases leads to significant losses in the aggregated DER flexibility potential. For example, as shown by the quantitative results in Table IV, a lack of flexible unit coordination leads to a flexibility area reduction of 15.72–29.23%, depending on phase. When voltage unbalance constraints (1o) are imposed, the reduction increases further to 34.80–60.28%, depending on phase. These observations suggest that at least 30% of theoretical aggregated flexibility may not be achievable in realistic unbalanced LV distribution networks.
Figure 9: Impact of phase coordination and voltage unbalance constraints on flexibility services in the real 221-bus system. The aggregated P-Q flexibility areas are shown for each phase. Cross markers indicate the initial operating point with no flexibility provision, while the coordinates represent the total network power consumption per phase. Flexibility services are provided by 12 single-phase units. Imposing phase coordination constraints prevents joint flexibility provision by units connected to different phases, thereby limiting voltage unbalance management and reducing the feasible area for each phase.
TABLE IV: Impact of Phase Coordination and Voltage Unbalance Constraints on Aggregated P-Q Flexibility in the 221-Bus System
Simulation Aggregated P-Q flexibility area per phase, kVAr2 (%)
Phase $\phi=\{a\}$ Phase $\phi=\{b\}$ Phase $\phi=\{c\}$
Perfect coordination of units, no voltage unbalance constraints $1721.72$  ( $100.00$ %) $1709.93$  ( $100.00$ %) $1702.88$  ( $100.00$ %)
No coordination of units, no voltage unbalance constraints $1211.63$  ( $70.37$ %) $1441.21$  ( $84.28$ %) $1300.69$  ( $76.38$ %)
No coordination of units, VUF limit of 1.0% imposed $1122.60$  ( $65.20$ %) $1050.17$  ( $61.42$ %) $676.34$  ( $39.72$ %)
No coordination of units, VUF limit of 0.9% imposed $1035.54$  ( $60.15$ %) $935.33$  ( $54.70$ %) $517.11$  ( $30.37$ %)
No coordination of units, VUF limit of 0.8% imposed $915.38$  ( $53.17$ %) $822.88$  ( $48.12$ %) $376.33$  ( $22.10$ %)
No coordination of units, VUF limit of 0.7% imposed $787.25$  ( $45.72$ %) $710.63$  ( $41.56$ %) $253.27$  ( $14.87$ %)
No coordination of units, VUF limit of 0.6% imposed $662.87$  ( $38.50$ %) $594.61$  ( $34.77$ %) $146.07$  ( $8.58$ %)
No coordination of units, VUF limit of 0.5% imposed $544.67$  ( $31.63$ %) $476.56$  ( $27.87$ %) $58.20$  ( $3.42$ %)
No coordination of units, VUF limit of 0.4% imposed $434.67$  ( $25.25$ %) $357.13$  ( $20.89$ %) $0.02$  ( $0.00$ %)
No coordination of units, VUF limit of 0.3% imposed $330.72$  ( $19.21$ %) $211.44$  ( $12.37$ %) $0.00$  ( $0.00$ %)
No coordination of units, VUF limit of 0.2% imposed $198.74$  ( $11.54$ %) $72.17$  ( $4.22$ %) $0.00$  ( $0.00$ %)
No coordination of units, VUF limit of 0.1% imposed $56.18$  ( $3.26$ %) $0.69$  ( $0.04$ %) $0.00$  ( $0.00$ %)
@@ IV-B Case Study: 221-bus LV Distribution Network in the UK
Note that in this simulation setting, all flexible units have been assigned the same P-Q capabilities and distributed evenly across phases. Yet, their operation is limited considerably due to phase coordination issues and voltage unbalance. It follows that the provision of flexibility services by DER can be challenging even in LV networks with low load unbalance as flexible units can exacerbate existing voltage unbalances by producing or consuming power. To effectively manage such unbalances and voltage constraints, units connected to different phases must be perfectly coordinated. The simulations performed for the realistic 221-bus network demonstrate that lack of coordination between units significantly reduces the potential of DER flexibility services and can magnify voltage unbalance problems. Other aspects and implications of these findings are discussed in Section V.
IV-C Stochastic Analysis
@@ V Discussion
The simulations performed in Section IV demonstrate the impacts of phase unbalance and imperfect coordination between distributed flexibility providers. Specifically, it is shown that the total DER flexibility potential can be significantly reduced due to voltage unbalance constraints and lack of coordination between flexible units connected to different phases. This highlights the importance of accurate nonlinear three-phase OPF models for the analysis of the network’s aggregated P-Q flexibility, as such models are capable of simulating voltage unbalances and coordination issues across phases. In this regard, the proposed flexibility quantification framework allows to directly translate voltage unbalance and phase coordination constraints as reductions of the aggregated P-Q flexibility. Thus, it can be used by DSO to realistically predict flexibility available in distribution networks and avoid severe voltage unbalances.
Yet, the proposed framework relies on several assumptions regarding the modelling of DER operation in unbalanced distribution networks. This section further discusses the implications of these assumptions, including the choice of DER placement locations, voltage unbalance definitions, DER coordination limitations, as well as the framework’s computational complexity and scalability.
V-A Impact of DER Location on Voltage Unbalance
@@
Multiple aspects of flexibility provision in unbalanced LV distribution networks have to be further investigated. One of the key factors affecting the aggregated P-Q flexibility is the relationship between the placement of flexible units and locations of voltage unbalances. For instance, in the 221-bus case study displayed in Fig. 7, the highest load unbalance is happening at the feeder with the flexible unit F5{a} connected. Six more flexible units are placed in this part of the network. Therefore, the design of this experiment inherently creates constraints on the operation of these flexible units due to the nearby voltage unbalances. That is, units must be coordinated when consuming or producing power to manage voltage constraints and keep the network’s operation feasible. For example, units connected to phase $\phi=\{a\}$ (the phase with the highest load) must not increase their power consumption while units in other phases consume power. Otherwise, voltage unbalance in some parts of the network may increase to unacceptable levels or even the voltage limits (1b) for phase $\{a\}$ may be violated. Note that this would not happen if flexible units were located far from the load unbalance, say, at the primary substation. Therefore, the question arises: how does the initial network unbalance impact the network’s P-Q flexibility and how does it restrict flexible resources at different locations? This question has implications for DER planning models. Depending on DSO metrics and objectives, some DER should be placed closer to the primary substation, while other flexible resources should be located closer to existing voltage unbalances to manage voltage constraints. A similar problem was considered in [Wang2021, Antic2024], where it was demonstrated that suboptimal selection of the DER connection phase can lead to voltage unbalance and limit DER capacity.
The provision of flexibility services can also be impacted considerably by the integration of EV charging stations. This impact can be especially pronounced in residential and workplace areas, where charging points are predominantly single-phase. In such scenarios, uncoordinated EV charging may reduce available flexibility due to tighter voltage and VUF constraints. Investigating EV-driven unbalance and its implications for flexibility aggregation is a valuable direction for future research.
Another important aspect of flexible resource placement is its impact on power losses. In the 221-bus system, the total active power losses amount to 2.4 kW (2.8%) for the initial operating point with no flexibility provision. The operation of flexible units can either reduce or exacerbate these losses, depending on their location and coordination. Therefore, investigating loss-minimising flexibility aggregation represents a promising direction for future work.
V-B Unbalance Definition and Modelling Assumptions
In this work, voltage unbalance in LV distribution networks is quantified and constrained using the VUF, defined as the ratio between the negative sequence and positive sequence of voltage components. This definition is widely adopted for electric power quality monitoring and is recommended by standards such as IEC 61000-2-2 and IEEE 1159 [nakadomari2024]. However, alternative definitions of voltage unbalance have been proposed in the literature. For example, voltage unbalance can be formulated as the ratio of the zero sequence to positive sequence. Voltage unbalance can also be quantified through the phase voltage unbalance rate (PVUR), computed based on the average phase voltage magnitude and the maximum deviation among phases. In addition, the current unbalance can be analysed using phase current magnitudes and current sequence components. It is therefore important for future research to systematically compare alternative unbalance definitions and analyse their impact on P-Q flexibility aggregation.
Another key modelling assumption made in this work is the use of a three-phase three-wire AC OPF formulation. This formulation captures the propagation of voltage unbalance along network lines and allows the explicit modelling of single-phase, two-phase, and three-phase load connections. However, it does not explicitly model the neutral conductor and grounding effects. Therefore, the proposed framework can be extended to four-wire network models. Analysis of the impact of the neutral conductor and grounding on aggregated P-Q flexibility is left for future research.
V-C Modelling of DER Coordination across Phases
A few modelling assumptions are made in this work regarding data availability, network observability, and DER coordination. Specifically, to demonstrate the fundamental impact of voltage unbalance constraints on flexibility aggregation, two operating conditions are compared: 1) a perfectly coordinated case, in which all DER are jointly optimised across phases, and 2) an uncoordinated case, in which DER connected to different phases are optimised independently, as imposed by constraint (1p). Both conditions are idealised. The “perfect coordination” assumption is not always valid in real LV distribution networks, as some units may not be observable or may experience communication delays. The “no coordination” assumption implies that units in other phases are inactive. However, in practice, uncoordinated units across different phases may continue to operate and inject or withdraw power. Moreover, only a subset of flexible units may be coordinated, while other units may be temporarily unavailable or uncoordinated. Therefore, future research should investigate the impact of partial DER coordination and communication constraints on P-Q flexibility aggregation.
V-D Computational Complexity and Scalability
@@
Regarding the computational aspects of flexibility quantification, the main computational burden comes from the need to iteratively solve the nonlinear optimisation problem (1a)-(1p). A single OPF problem with no voltage unbalance constraints converges in 0.1-0.2 seconds and 5-50 seconds for the 5-bus and 221-bus cases, respectively. This means that the boundary of an aggregated P-Q flexibility area with 50 points can be estimated in less than 2500 seconds (43 minutes). On average, it takes 16 minutes to calculate a single boundary presented in Fig. 9. Note that these values are reported for the simulations performed on a laptop with an Intel Core i7-10510U CPU 1.80GHz and 16 GB of RAM. The flexibility estimation process can be greatly sped up by using more powerful computers and parallel computing (solving OPF problems for different points on different cores and machines). Therefore, DSO can perform this flexibility assessment close to real time. This implies that such flexibility areas can be updated during intraday markets, e.g., every 15 minutes. Yet, it is important to highlight that the proposed framework is not intended for real-time control. Instead, it is best suited for operational planning and market-related applications.